def pgcd(a,b) :
    while a%b != 0:
        a, b = b, a%b
    return b
print('Le PGCD de 57 et 21 est : ',pgcd(57,21))

Console

Le PGCD de 57 et 21 est : 3

Un text à l’envers

\newcommand{\reverse}[1]{\py{"#1"[::-1]}}

tset nu tse iceC